Subject: APO 3.5m users committee minutes, June 2012
From: Suzanne Hawley
Submitted: Wed, 22 Aug 2012 09:11:14 -0700 (PDT)
Message number: 1285
(previous: 1284,
next: 1286
up: Index)
APO 3.5-m Users Committee Phonecon, 6/11/12
Attending: Suzanne Hawley, Mark Klaene, Al Harper, John Stocke (for John
Bally), Scott Anderson, Bill Ketzeback, Sean Moran, Jon Holtzman, Rene
Walterbos, Jenny Greene (Princeton rep)
Absent: Remy Indebetouw, Bruce Gillespie
**********************************
User feedback and comments from institutional representatives:
o Chicago (Al Harper) - Al had nothing new to report.
o Washington (Scott Anderson) - Scott had nothing new to report.
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Princeton (Ed Turner) - Jenny had nothing new to report.
o Colorado (John Stocke) - comments given below during meeting.
o Virginia (Remy Indebetouw) - No report.
o Johns Hopkins (Sean Moran) - Sean had nothing new to report.
**********************************
Discussion of telescope/instruments report:
Mark gave an overview of the weather, telescope, and instrument report.
In addition, he mentioned the Little Bear fire in Ruidoso and smoke from
the Gila fires.
There was some discussion on the NIC-FPS status report. John Stocke
mentioned that Ed Chang from Goddard has offered help. The instrument
would have to go to Goddard. John Stocke will investigate the cost and
scope of work but there was mention that it would be low-cost compared to
having UVa work on the corrupted-image problem. A conversation on this
topic has started between Ed Chang and Jim Greene, and Stocke will
investigate further and report details back to Mark and Suzanne.
There was additional discussion regarding the NIC-FPS etalon. Etalon work
could still proceed at UVa. CU is open to Fred Hearty and the UVa
instrument group working on it. Waiting to hear back from UVa on their
interest in taking this on.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
05/03/12 through 05/30/12
0) Overview
Telescope continued to run well, but the overall good weather was marred
by smoke from a large fire in the Gila National Forest to our west. The
observers have been doing well, closing only when smoke levels rise to
extreme conditions and using the louvers and pressurization fan when
possible to keep the shutter open. The local fire danger dropped to
moderate with arrival of some rain this month.
1) Telescope
The telescope has been operational with no issues of note.
2) Instruments
DIS Blue still shows some signs of contamination after vacuum servicing.
Significant dewar work is planned for the summer and we are focusing on
keeping it operational until then.
Agile had no issues this period.
SPIcam is still running 5-10 Deg K warmer than we would expect with a good
vacuum and new coldhead. Average temperature is around 180K and it varies
with ambient temperature -- warmer during the day, cooler at night. Dark
current appears to be very good still at these temperatures. The camera
is ready for use. We are continuing to investigate the temperature issue.
The TSpec maskfile has been remade and tested at various binning factors.
Some issues have been isolated and we are looking at other ways to make
improvements.
Echelle is operational. Plan is for installation of the new grating
during summer shutdown. The new ECam had a shutter failure due to a
failed motherboard. Ed and Joe built an external driver and Fritz
implemented a software change that got Echelle operational again over the
Memorial Day weekend -- thanks very much for their efforts. A permanent
repair to the motherboard is in work with the manufacturer.
NIC-FPS has been operating in shared-risk observing mode due to occasional
image corruption issues. After last meeting's discussion, we are
currently discussing with UVa a proposal to address the etalon issue and
evaluate the possibility of an upgrade path.
**********************************
Summer meetings, scheduling
Suzanne will be out of the office this summer, so she suggested skipping
July and August. The tentative next meeting will be Monday Sept 10.
Preparation for the annual ARC board meeting will be earlier this year.
Please tell your institutional members to start preparing science slides
by the end of August so you can get them to Suzanne and Mark in September.
Time-trade requests while Suzanne is out will be handled primarily by
Russet McMillan, with oversight by Mark Klaene and Rene Walterbos.
Please send ToO and time-trade requests to all of Suzanne, Russet, Mark
and Rene (email addresses in cc line above).
A reminder that summer shutdown for the 3.5-m starts on July 23 with a
return to science on August 10th. Echelle may not return to science until
August 20th to give extra time for alignment and characterization of
its new grating.
**********************************
New Observing Specialist position:
Mark said that we are looking for a replacement for Gabrelle Saurage, who
will be leaving us this summer. The position requires a B.S. at a
minimum. There is a posting on the APO website as well as in the AAS Job
Register.
** note added in August - new Observing Specialist Alysha Shugart has been
hired. Gabrelle's last day was August 13. We appreciate all her effort
and hard work for the Observatory in the last several years.
**********************************
New DIS slitviewer camera:
Mark queried the Users Committee members about increasing the FOV to 4
arcminutes. Bill Ketzeback has been looking into possible designs. John
Stocke thought this sounded good. No one else had comments. Bill will be
looking into this again after the shutdown.
**********************************
Other business:
Jon Holtzman brought up an issue regarding how to handle alternative
scheduling and backup programs when a scheduled instrument is unavailable.
He asked if it was advisable to start an observatory-wide backup program,
most likely as service observing. John Stocke wanted to know if this
should be done more informally on an institution-by-institution basis, by
giving each observer a cache of targets with backup instruments. Suzanne
suggested we talk more about this in the future.
John Stocke added that CU would like to test a commercial off-the-shelf
tip-tilt mechanism on the 3.5-m. John Bally and he would like to
investigate this further. This would be basically applying an
amateur-type system to the eyepiece mount for the 3.5-m with a CCD camera
as a study for a system we could incorporate for imagers such as NIC-FPS
and SPIcam. Jon Holtzman also expressed an interest in trying such tests
on the NMSU 1-m. Mark and Jon Holtzman should be contacted for details.
Suzanne suggested such tests could be scheduled for Quarter 4.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Send new ARC partner suggestions to Suzanne and Rene (before
next meeting).
Status: Suzanne has received a small number of suggestions, and is still
looking for more. Please send your suggestions of an institution and a
contact person to her and Rene. Within two years Princeton will be
leaving ARC, and Chicago will also be looking at parting ways. This
leaves 1/3 of 3.5-m time available soon. We are pursuing a targeted
approach for now. A few feelers have produced some positive feedback.
[new actions from this meeting]:
===> ACTION: Science slides for BOG meeting should be to Suzanne and Mark
by September.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday
10 September at 8:30 AM Pacific Time. The agenda and other materials
will be sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1285 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO